# Fate of the Unbound States in Near-infinitely Deep Potential Models

### 23 Apr 2026 v1

@@ Abstract
Based on the one-dimensional model with quasi-periodicity and nearly infinite depth potential well, this paper studies how the depth of the potential well and non-Hermiticity affects the unbound states. By extending the Liu-Xia model to a deeper structure, we confirm through calculating IPR and based on Avila’s global theory, that although the potential well is deeper, there are still unbound states within specific energy intervals. Extending the research to non-Hermitian systems with gain-loss effects, we find that the non-Hermiticity leads to the existence of unbound states in a mixed state form composed of bound states and unbounded states. However, there are clear energy boundaries between the mixed regions with unbound states and the pure bound state regions, which can be proved by Avila’s global theory. Our research results provide new insights into the unbound states in extreme potential fields.
@@ I Introduction
In the framework of quantum mechanics, the eigenstate properties of particles in potential wells have always been a core research topic, and the classification of bound and unbound states is the basis for understanding quantum transport and metallic-insulating transition Shankar (1994). For the one-dimensional infinite deep potential well, the textbook conclusion holds that the particle can only exist in discrete bound states, and the unbound scattering states can only be realized in the finite deep or bounded potential field Stillinger and Herrick (1975); Marinica et al. (2008). The traditional definition divides the quantum states by the particle’s eigenvalue relative to the potential barrier height: when the eigenvalue is lower than the potential barrier, the particle is confined in the potential well to form a bound state; when the eigenvalue is higher than the potential barrier, the particle escapes the potential well to form an unbound state Capasso et al. (1992), which is accompanied by the correspondence between discrete spectrum and bound state, continuous spectrum and unbound state Molina et al. (2012); Gorkunov et al. (2020).
@@
However, the discovery of bound states in the continuous spectrum by von Neumann and Wigner in 1929 broke the inherent cognition of the correspondence between quantum states and energy spectra von Neumann and Wigner (1929), and a series of follow-up studies Pursey and Weber (1995); Weber and Pursey (1998) have found similar exotic quantum states in various potential field structures, which indicates the richness and complexity of quantum state characteristics in non-traditional potential systems Koshelev et al. (2018); Kang et al. (2021). Until recently, Liu and Xia proposed a one-dimensional discrete quasiperiodic near-infinitely deep potential model $V(n)=V\sec^{2}(2\pi\alpha n)$ ($V$ is the potential strength), which for the first time proved that unbound states can exist in the potential field that tends to near-infinitely depth Liu and Xia (2023). This model is different from the traditional infinite deep potential well in finite space: its potential strength approaches infinity only when the lattice number $n\to\infty$, and the critical energy $E=2t$ ($t$ is the unit of energy) is the mobility edge that separates bound and unbound states. When $E<2t$, the quantum state is in an unbound state that spreads over the entire system space; when $E>2t$, the quantum state forms a localized bound state Liu and Xia (2023); Dwiputra and Zen (2022). This research breaks out the classical understanding of infinite deep potential wells and opens up a new research direction for the study of unbound states in extreme potential fields.
@@
On the one hand, in traditional potential energy systems, the depth of the potential well is a key parameter governing the quantum states of particles: for a finite-depth potential well, an increase in potential depth enhances the localization of particles, reduces the tunneling probability of the wave function, and even converts unbound states into bound states Shankar (1994). For a near-infinitely deep potential well, it is known that unbound states are confined to the spectral region of $E<2t$ Liu and Xia (2023), yet how an increase in potential depth affects unbound states remains an unsolved problem. On the other hand, with the development of non-Hermitian quantum mechanics, the combination of non-Hermitian effects and Anderson localization has been found to induce a series of exotic quantum phenomena, such as non-Hermitian Anderson localization Longhi (2019); Jiang et al. (2019), non-Hermitian mobility edges Liu et al. (2020a, b); Xu et al. (2021); Peng et al. (2023); Lin et al. (2022); Jiang et al. (2025), non-Hermitian mobility edge rings Li and Li (2024). In the Hermitian system with a near-infinitely deep potential well, the mobility edge is known to be a real spectral line that separates bound states from unbound states Liu and Xia (2023). However, it remains an open question whether the introducing of non-Hermitian effects will induce diverse spectral features and modify the fate of unbound states. Based on these considerations, in this study, we take the quasiperiodic near-infinitely deep potential model as the research object and systematically investigate the effects of enhanced potential well depth and non-Hermitian effects on the unbound states of the system.
This paper is organized as follows. Section. II studies the unbound states under deeper potential well. Section. III studies the unbound states under non-Hermitian effects. The summary and outlook are presented in Sec. IV.
@@
Figure 1: (Color online) The sketch of the near-infinitely deep potential wells (a) $V_{1}(n)/V=\frac{1}{\cos(2\pi\alpha n)^{2}}$. (b) $V_{2}(n)/V=\frac{1+\sin(2\pi\alpha n)^{2}}{\cos(2\pi\alpha n)^{2}}$. It is seen that the potential well $V_{2}$ is much deeper than $V_{1}$.
@@ II The deeper potential model
Liu and Xia studied a near-infinitely potential model with on-site potential $V_{1}(n)/V=\sec(2\pi\alpha n)^{2}$ ($V$ is the tunable parameter of potential and $\alpha=\frac{\sqrt{5}-1}{2}$ Liu and Xia (2023). At lattice site $n$, the magnitude of the potential is much greater than the order of $10^{4}$ (see the sketch in Fig. 1(a)), which gives rise to the term ”near-infinitely deep potential” as used. Considering the particle hops between the nearest-neighboring lattice sites, it was found that there were unbound states in the energy domain $E<2t$ ($t$ is the energy unit). In the following, we are motivated to study the fate of unbound state under deeper potential well situation. The deeper potential is described by $V_{2}(n)/V=\frac{1+\sin(2\pi\alpha n)^{2}}{\cos(2\pi\alpha n)^{2}}$, and the sketch of the deeper potential is presented in Fig. 1(b). It is seen that after considering the additional factors, the depth of the potential well has significantly increased.
@@
$\hat{H}=\sum_{n}t\hat{c}^{n+1}\hat{c}_{n}+H.c.+\sum_{n}V_{2}(n)\hat{c}_{n}\hat{c}_{n},$ (1)
with $V_{(}n)=V\frac{1+\sin(2\pi\alpha n)^{2}}{\cos(2\pi\alpha n)^{2}}$.
Figure 2 presents the unbound-bound phase diagram when the system size is $L=1597$. The colors represent the magnitude of the inverse participation ratio (IPR). The IPR of the $j$-th normalized wave function $|\psi^{j}\rangle=\sum_{n}\phi(n)\hat{c}_{n}^{{\dagger}}|0\rangle$ is defined as ${\rm IPR}=\sum_{n}|\phi(n)|^{4}$ Biddle and Das Sarma (2010). For unbound states, if the wave function is extended, IPR is proportional to $1/L$, which tends to zero as the system size tends to infinity. For bound states, if the wave function is localized, IPR is a finite-sized number Liu and Xia (2023); Ganeshan et al. (2015); Liu et al. (2022). As is evident from the figure, increasing the potential well depth only reduces the energy range of the bound states without causing their disappearance. The unbound states occur in the energy domain where $E_{c2}<E<E_{c1}$ with $E_{c1}=2t-V$ and $E_{c2}=-2t-V$, while the bound states occur in the energy domain where $E>E_{c1}$. To visualize the characteristics of unbound states, Fig. 3 shows the distributions of wave functions for four distinct energies at potential parameter $V=0.5t$, with energies located on either side of the critical energy $E_{c1}$. As seen from Fig. 3, for $E<E_{c1}$ (Figs. 3(a), 3(b)), the wave functions are extended across the entire space, corresponding to unbound states. In contrast, for $E>E_{c1}$ (Figs. 3(c), 3(d)), the wave functions are strongly localized in finite spatial regions and decay to zero at the boundaries, indicative of bound states. The results are in agreement with the IPR analysis.
Figure 2: (Color online) The energies $E$ versus $V$ of the deeper potential model. The colors denote the values of the IPR. The two blue solid lines denote the critical energies $E_{c1}=2t-V$ and $E_{c2}=-2t-V$, respectively. Within the energy domain $\left[E_{c2},E_{c1}\right]$, there are unbound states. Above $E_{c1}$, there are bound states. The system size is $L=1597$.
Figure 3: (Color online) Four wave funcitons of the deeper potential model. (a) The unbound state at E=0.5593tE=-0.5593t; (b) The unbound state at E=0.5568tE=-0.5568t; (c) The bound state at $E=3.1698t$; (d) The bound state at $E=12.8646t$. The system size is $L=1597$.
We argue that the unbound-bound transition can be exactly proved by analyzing the Lyapunov exponent (LE) $\gamma$. For unbound state, $\gamma=0$; while for bound state, $\gamma>0$. The $\gamma$ can be determined from the product of the transfer matrices. With wave function $|\psi^{j}\rangle=\sum_{n}\phi(n)\hat{c}_{n}^{{\dagger}}|0\rangle$ and energy $E$, the Schrödinger equation of amplitudes $\phi_{j}$ in transfer matrix form is written as
@@
To calculate $\gamma^{B}(E)$, one should employ the Avila’s global theory Avila (2015). The procedure starts with an analytical continuation of the phase, namely, $\theta\rightarrow\theta+i\epsilon$. As $\epsilon\rightarrow\infty$, the direct evaluation of $B_{n}$ leads to
@@
Then we obtain $\gamma^{B}(E)=2\epsilon-2\ln 2+\max\{\ln|\frac{\frac{E+V}{t}\pm\sqrt{\left(\frac{E+V}{t}\right)^{2}-4}}{2}|\}$. According to Avila’s global theory, $\gamma^{B}(E)$ is a convex and piecewise linear function with integer slope. It can be concluded that the Lyapunov exponent can be determined when $\epsilon$ returns to 0. Therefore, the LE of a wave function is
@@
From the above equation, we know that when $|E+V|>2t$, $\gamma(E)>0$, while $|E+V|<2t$, $\gamma(E)\equiv 0$. Therefore, the bound states exist in the energy domain $E>2t-V\bigcup E<-2t-V$ and the unbound states locate in the energy domain 2tV<E<2tV-2t-V<E<2t-V. In fact, from the IPR results, we know that below E=2tVE=2t-V, there have been already unbound states. Thus, in this deeper potential model, E=2tVE=2t-V can be regarded as the zero potential energy surface.
@@ III Non-Hermitian effect on unbound states
The intriguing discovery of unbound states in a near-infinitely deep potential well challenges the conventional understanding of wave function behavior in deep potential systems, uncovering counterintuitive quantum dynamical properties that contradict the descriptions of infinite deep potential well physics in standard textbooks. The aforementioned work has systematically elucidated the existence and characterization rules of unbound states in the Hermitian quasiperiodic potential model. Non-Hermiticity, a ubiquitous feature in realistic quantum systems, arises from the coupling between the system and its environment as well as gain and loss effects, and its incorporation induces novel quantum phenomena such as the modulation of real-complex energy spectra and delocalization-localization properties. Building on this foundation, we extend the research framework of the near-infinitely deep and deeper potentials to the non-Hermitian regime, aiming to investigate the regulatory laws governing how non-Hermitian effects modulate the existence forms and spatial distributions of unbound states.
We firstly investigate the effects of non-Hermiticity on the Liu-Xia model. Induced by on-site gain and dissipation effects, the non-Hermitian potential is expressed as $V^{\rm NH}_{1}(n)=\frac{V}{\cos(2\pi\alpha n+\theta+ih)^{2}}\equiv V\sec(2\pi\alpha n+\theta+ih)^{2}$. The Schröedinger equation is then expressed in the following form:
@@
Then, the LE of non-Hermitian system, namely $\gamma^{\rm NH1}(E)$, is expressed as
@@
To calculate $\gamma^{B}_{\rm NH1}(E)$, we still employ the Avila’s global theory. After performing the analytic continuation on $B^{\rm NH1}_{n}$, that is $\theta\rightarrow\infty$, we have
@@
Then the $\gamma^{B}_{\rm NH1}(E)$ is given by
@@
According to the above Eq. (20), we know that when $|E|>2t$, $\gamma^{\rm NH1}(E)>0$, while $|E|<2t$, $\gamma^{\rm NH1}(E)\equiv 0$. This result is highly similar to that of the Liu-Xia model, except that the energy $E$ here refers to the real part of the complex energy spectrum, which will be manifested in the subsequent IPR results. The analytical results predict that the unbound states will exist in the real energy domain 2t<E<2t-2t<E<2t, while the bound states will exist in the real energy domain $E>2t\bigcup E<-2t$. For convenience, we mark $E^{\rm NH1}_{c1}=2t$ and $E^{\rm NH1}_{c2}=-2t$. The label $E$ in the following denotes the real part energy.
@@
Figure 4: (Color online) The energies $E$ versus $V$ of the non-Hermitian Liu-Xia model. The colors denote the values of the IPR. The two blue solid lines denote the critical energies $E^{\rm NH1}_{c1}=2t$ and $E^{\rm NH1}_{c2}=-2t$, respectively. The energy domain $\left[E_{c2},E_{c1}\right]$ is a mixed region where unbound states and bound states coexist. Above $E_{c1}$ and below $E_{c2}$, there are bound states. The system size is $L=1597$.
@@
Figure 5: (Color online) Four wave functions of the non-Hermitian Liu-Xia model. (a) The unbound state at E=45.7662tE=-45.7662t; (b) The unbound state at E=0.7628tE=-0.7628t; (c) The bound state at E=0.7625tE=-0.7625t; (d) The bound state at $E=1.1534t$. The involved parameters are $h=0.1$,$V=0.5t$, and $L=1597$.
@@
In Fig. 4, we present the real part of energies $E$ (arranged in ascending order) of the non-Hermitian Liu-Xia model as a function of the potential parameter $V$. The color coding indicates the magnitude of the IPR, and the blue lines mark the critical energies $E^{\rm NH1}_{c1}$ and $E^{\rm NH1}_{c2}$ obtained from analytical derivations. Similar to the Hermitian scenario, the energy region $E>E^{\rm NH1}_{c1}$ still accommodates bound states. Differently, bound states emerge in the lower real energy domain with $E<E_{c2}$. Furthermore, it can be seen that unbound states do not vanish due to non-Hermiticity, but remain present in the energy window $E^{\rm NH1}_{c2}<E<E^{\rm NH1}_{c1}$. Different from the Hermitian systems, the conventional non-Hermitian mobility-edge and intermediate phase systems, where unbound states fully occupy one or several entire energy domains Liu and Xia (2023); Liu et al. (2020a); Xu et al. (2021); Peng et al. (2023); Biddle and Das Sarma (2010); Ganeshan et al. (2015); Liu et al. (2022), unbound states here appear in the form of mixed phase consisting of both unbound and bound states without a regular pattern. It is readily seen that the real part energies of unbound states are embedded in the real part energies of bound states. The emergent mixed phase indicates that non-Hermitian nature can fundamentally reconstruct the energy spectrum and localization structure of the system, leading to uncommon phenomenon.
@@
Figures 5(a)-5(d) shows the spatial distribution of the wave functions under four different energies. In Fig. 5(a), the real part of the energy corresponding to the wave function is located below the critical energy $E^{\rm NH}_{c1}$. It can be observed that its spatial distribution state is localized and it is a bound state. The real parts of the wave functions in Figs. 5(b)-5(c) are located within the energy domain $E^{\rm NH1}_{c2}<E<E^{\rm NH1}_{c1}$. The distributions in Figs. 5(b) and 5(c) are extended and unbound states, while the distribution in Fig. 5(d) is localized and is a bound state. This result clearly demonstrates that in non-Hermitian cases, within the energy domain $E^{\rm NH1}_{c2}<E<E^{\rm NH1}_{c1}$, bound states and unbound states coexist, which is consistent with the results of IPR. This also indicates that the analytical solution derived based on Avila’s global theory can predict both the energy domain of unbound states and the energy domain of pure bound states.
@@
Next, we study the non-Hermitian effect on the deeper potential model. Considering the on-site gain and loss on $V_{2}(n)$, the resulting non-Hermitian potential $V^{\rm NH2}_{2}(n)$ is given by
$V^{\rm NH2}_{1}(n)$ $=\frac{V\left[1+\sin^{2}(2\pi\alpha n+\theta+ih)\right]}{\cos^{2}(2\pi\alpha n+\theta+ih)}$ (21)
@@
Following the previous strategy, here we still express $T^{\rm NH2}_{n}$ as the product of two parts, i.e., $T^{\rm NH2}_{n}=A^{\rm NH2}_{n}B^{\rm NH2}_{n}$, so we have
@@
It is noted that $A^{\rm NH1}_{n}=A^{\rm NH2}_{n}$, therefore we have
@@
To obtain $\gamma^{B}_{\rm NH2}$, we again employ Avila’s global theory. After performing analytic continuation ($\theta\rightarrow\theta+i\epsilon$) on the matrix $B^{\rm NH2}_{n}$, its expression in the infinite-$\epsilon$ limit reads
@@
Then, the $\gamma^{B}_{\rm NH2}$ is given by

### 27 Apr 2026 v2

# Unbound States and Mixed Bound–Unbound Phases in Near-Infinitely Deep Potentials

@@ Abstract
We investigate the robustness of unbound states in one-dimensional quasiperiodic models with near-infinitely deep potentials. By constructing a deeper extension of the Liu-Xia model and combining inverse participation ratio (IPR) calculations with Lyapunov-exponent analysis based on Avila’s global theory, we show that increasing the potential depth does not eliminate unbound states. Instead, it shifts and narrows their energy window to 2tV<E<2tV-2t-V<E<2t-V. We further extend the analysis to non-Hermitian quasiperiodic potentials with gain and loss. In these systems, unbound states survive within analytically determined real-energy intervals, but they no longer occupy the whole interval uniformly; rather, they coexist with bound states and form a mixed bound-unbound phase. The corresponding boundaries between the mixed region and the pure bound-state regions are obtained exactly from the Lyapunov exponent. These results demonstrate that unbound states in extreme quasiperiodic potentials are controlled not only by the potential depth but also by the spectral and localization structures induced by non-Hermiticity.
@@ I Introduction
In the framework of quantum mechanics, the eigenstate properties of particles in potential wells have always been a core research topic, and the classification of bound and unbound states is the basis for understanding quantum transport and metal-insulator transitions [22]. For a one-dimensional infinitely deep potential well, the textbook conclusion holds that the particle can only exist in discrete bound states, and the unbound scattering states can only be realized in a finite-depth or bounded potential field [23, 18]. The traditional definition divides the quantum states by the particle’s eigenvalue relative to the potential barrier height: when the eigenvalue is lower than the potential barrier, the particle is confined in the potential well to form a bound state; when the eigenvalue is higher than the potential barrier, the particle escapes the potential well to form an unbound state [3], which is accompanied by the correspondence between discrete spectrum and bound state, continuous spectrum and unbound state [19, 6].
@@
However, the discovery of bound states in the continuous spectrum by von Neumann and Wigner in 1929 broke the inherent cognition of the correspondence between quantum states and energy spectra [24], and a series of follow-up studies [21, 25] have found similar exotic quantum states in various potential-field structures, which indicates the richness and complexity of quantum state characteristics in non-traditional potential systems [10, 9]. Until recently, Liu and Xia proposed a one-dimensional discrete quasiperiodic near-infinitely deep potential model $V(n)=V\sec^{2}(2\pi\alpha n)$ ($V$ is the potential strength), which for the first time proved that unbound states can exist in the potential field that tends to near-infinite depth [13]. This model is different from the traditional infinite deep potential well in finite space: its potential strength approaches infinity only when the lattice site $n\to\infty$, and the critical energy $E=2t$ ($t$ is the unit of energy) is the mobility edge that separates bound and unbound states. When $E<2t$, the quantum state is in an unbound state that spreads over the entire system space; when $E>2t$, the quantum state forms a localized bound state [13, 4]. This research breaks with the classical understanding of infinite deep potential wells and opens up a new research direction for the study of unbound states in extreme potential fields.
@@
On the one hand, in traditional potential energy systems, the depth of the potential well is a key parameter governing the quantum states of particles: for a finite-depth potential well, an increase in potential depth enhances the localization of particles, reduces the tunneling probability of the wave function, and even converts unbound states into bound states [22]. For a near-infinitely deep potential well, it is known that unbound states are confined to the spectral region of $E<2t$ [13], yet how an increase in potential depth affects unbound states remains an unsolved problem. On the other hand, with the development of non-Hermitian quantum mechanics, the combination of non-Hermitian effects and Anderson localization has been found to induce a series of exotic quantum phenomena, such as non-Hermitian Anderson localization [17, 7], non-Hermitian mobility edges [14, 16, 26, 20, 12, 8], non-Hermitian mobility edge rings [11]. In the Hermitian system with a near-infinitely deep potential well, the mobility edge is known to be a real spectral line that separates bound states from unbound states [13]. However, it remains an open question whether the introduction of non-Hermitian effects will induce diverse spectral features and modify the fate of unbound states. Based on these considerations, in this study, we take the quasiperiodic near-infinitely deep potential model as the research object and systematically investigate the effects of enhanced potential well depth and non-Hermitian effects on the unbound states of the system. In the present lattice quasiperiodic models, we identify unbound states with spatially extended eigenstates characterized by a vanishing IPR and a zero Lyapunov exponent, whereas bound states correspond to localized eigenstates with a finite IPR and a positive Lyapunov exponent. This identification allows us to address two central questions: whether unbound states survive when the near-infinitely deep potential is further deepened, and whether non-Hermitian gain and loss destroy or reshape the corresponding unbound-state regions. We show that unbound states persist in the deeper potential model, although their energy window is shifted to 2tV<E<2tV-2t-V<E<2t-V. In the non-Hermitian extensions, unbound states survive but appear in mixed bound-unbound phases within analytically determined real-energy intervals.
The rest of this paper is organized as follows. Section. II analyzes the fate of unbound states in a deeper near-infinitely deep potential. Section. III extends the discussion to non-Hermitian potentials and identifies the resulting mixed bound-unbound phases. A summary is given in Sec. IV.
@@
Figure 1: (Color online) The sketch of the near-infinitely deep potential wells (a) $V_{1}(n)/V=\sec^{2}(2\pi\alpha n)$. (b) $V_{2}(n)/V=\frac{1+\sin(2\pi\alpha n)^{2}}{\cos(2\pi\alpha n)^{2}}$. The comparison illustrates how the modified potential enhances the depth while preserving the quasiperiodic near-singular structure.
@@ II The deeper potential model
Liu and Xia studied a one-dimensional quasiperiodic model with the near-infinitely deep on-site potential $V_{1}(n)/V=\sec^{2}(2\pi\alpha n)$, where $V$ is the tunable potential strength and $\alpha=(\sqrt{5}-1)/2$ [13]. As illustrated in Fig. 1(a), the potential can reach extremely large values at certain lattice sites, which motivates the term near-infinitely deep potential. For nearest-neighbor hopping, this model supports unbound states in the energy region $E<2t$, with $t$ taken as the energy unit. We now ask whether these unbound states survive when the potential is made even deeper. To this end, we introduce the modified potential $V_{2}(n)/V=\frac{1+\sin(2\pi\alpha n)^{2}}{\cos(2\pi\alpha n)^{2}}$, whose profile is shown in Fig. 1(b). The additional numerator enhances the potential depth while retaining the same quasiperiodic near-singular structure.
@@
$\hat{H}=\sum_{n}t\hat{c}_{n+1}^{\dagger}\hat{c}_{n}+H.c.+\sum_{n}V_{2}(n)\hat{c}_{n}^{\dagger}\hat{c}_{n},$ (1)
with $V_{2}(n)=V\frac{1+\sin^{2}(2\pi\alpha n)}{\cos^{2}(2\pi\alpha n)}$.
Figure 2 presents the unbound-bound phase diagram when the system size is $L=1597$. The colors represent the magnitude of the inverse participation ratio (IPR). The IPR of the $j$-th normalized wave function $|\psi^{j}\rangle=\sum_{n}\phi(n)\hat{c}_{n}^{{\dagger}}|0\rangle$ is defined as ${\rm IPR}=\sum_{n}|\phi(n)|^{4}$ [2]. This quantity provides a direct localization diagnostic: for an extended unbound state, ${\rm IPR}\sim 1/L$ and therefore vanishes in the thermodynamic limit, whereas for a localized bound state, the IPR remains finite [13, 5, 15]. As is evident from the figure, increasing the potential well depth only reduces the energy range of the unbound states without causing their disappearance. The unbound states occur in the energy domain $E_{c2}<E<E_{c1}$ with $E_{c1}=2t-V$ and $E_{c2}=-2t-V$, while the bound states occur in the energy domains $E>E_{c1}$ and $E<E_{c2}$. To visualize the characteristics of unbound states, Fig. 3 shows the distributions of wave functions for four distinct energies at potential parameter $V=0.5t$, with energies located on either side of the critical energy $E_{c1}$. As seen from Fig. 3, for $E<E_{c1}$ (Figs. 3(a), 3(b)), the wave functions are extended across the entire space, corresponding to unbound states. In contrast, for $E>E_{c1}$ (Figs. 3(c), 3(d)), the wave functions are strongly localized in finite spatial regions and decay to zero at the boundaries, indicative of bound states. The results are in agreement with the IPR analysis. A further finite-size scaling analysis of representative states would provide an additional confirmation that the low-IPR states obey the expected ${\rm IPR}\sim 1/L$ behavior in the thermodynamic limit.
Figure 2: (Color online) The energies $E$ versus $V$ of the deeper potential model. The colors denote the values of the IPR. The two blue solid lines denote the critical energies $E_{c1}=2t-V$ and $E_{c2}=-2t-V$, respectively. Within the energy domain $\left[E_{c2},E_{c1}\right]$, there are unbound states. Above $E_{c1}$, there are bound states. The persistence of low-IPR states inside this interval demonstrates that increasing the potential depth does not eliminate unbound states. The system size is $L=1597$.
Figure 3: (Color online) Four wave functions of the deeper potential model. (a) The unbound state at E=0.5593tE=-0.5593t; (b) The unbound state at E=0.5568tE=-0.5568t; (c) The bound state at $E=3.1698t$; (d) The bound state at $E=12.8646t$. The system size is $L=1597$.
We argue that the unbound-bound transition can be exactly proved by analyzing the Lyapunov exponent (LE) $\gamma$. For an unbound state, $\gamma=0$, while for a bound state, $\gamma>0$. The $\gamma$ can be determined from the product of the transfer matrices. With wave function $|\psi^{j}\rangle=\sum_{n}\phi(n)\hat{c}_{n}^{{\dagger}}|0\rangle$ and energy $E$, the Schrödinger equation of amplitudes $\phi_{j}$ in transfer matrix form is written as
@@
To calculate $\gamma^{B}(E)$, one should employ Avila’s global theory [1]. The procedure starts with an analytical continuation of the phase, namely, $\theta\rightarrow\theta+i\epsilon$. As $\epsilon\rightarrow\infty$, the direct evaluation of $B_{n}$ leads to
@@
Then, we obtain $\gamma^{B}(E)=2\epsilon-2\ln 2+\max\{\ln|\frac{\frac{E+V}{t}\pm\sqrt{\left(\frac{E+V}{t}\right)^{2}-4}}{2}|\}$. According to Avila’s global theory, $\gamma^{B}(E)$ is a convex and piecewise linear function with integer slope. The quantization of the acceleration, together with this convex piecewise-linear structure, fixes the physical Lyapunov exponent once the large-$\epsilon$ asymptotic form is known. It can be concluded that the Lyapunov exponent can be determined when $\epsilon$ returns to 0. Therefore, the LE of a wave function is
@@
From the above equation, we know that when $|E+V|>2t$, $\gamma(E)>0$, while when $|E+V|<2t$, $\gamma(E)\equiv 0$. Therefore, the bound states exist in the energy domains E>2tVE>2t-V and E<2tVE<-2t-V and the unbound states are located in the energy domain 2tV<E<2tV-2t-V<E<2t-V. In fact, from the IPR results, we know that below E=2tVE=2t-V, there are already unbound states. Thus, in this deeper potential model, E=2tVE=2t-V can be regarded as the zero potential energy surface.
@@ III Non-Hermitian effect on unbound states
The existence of unbound states in near-infinitely deep quasiperiodic potentials challenges the conventional expectation that an increasingly deep potential necessarily enhances confinement. The Hermitian analysis above shows that the fate of these states is governed by analytically determined Lyapunov-exponent boundaries rather than by the potential depth alone. We now extend this framework to non-Hermitian systems, where gain and loss can generate complex spectra and modify localization properties. The central question is whether such non-Hermitian effects destroy the unbound states or reorganize them into new spectral-localization structures.
We first investigate the effects of non-Hermiticity on the Liu-Xia model. With on-site gain and dissipation effects, the non-Hermitian potential is expressed as $V^{\rm NH}_{1}(n)=\frac{V}{\cos(2\pi\alpha n+\theta+ih)^{2}}\equiv V\sec(2\pi\alpha n+\theta+ih)^{2}$. For the non-Hermitian models considered below, the IPR is calculated from normalized right eigenstates. Since the eigenenergy is generally complex, we write it as $\mathcal{E}=E+i\eta$ and use $E\equiv{\rm Re}(\mathcal{E})$ to denote the real-part energy unless otherwise specified. The Schrödinger equation is then expressed in the following form:
@@
Then, the LE of the non-Hermitian system, namely $\gamma^{\rm NH1}(E)$, is expressed as
@@
To calculate $\gamma^{B}_{\rm NH1}(E)$, we still employ Avila’s global theory. After performing the analytic continuation on $B^{\rm NH1}_{n}$, that is, $\theta\rightarrow\theta+i\epsilon$, we have
@@
Then, $\gamma^{B}_{\rm NH1}(E)$ is given by
@@
According to Eq. (20), we know that when $|E|>2t$, $\gamma^{\rm NH1}(E)>0$, while $|E|<2t$, $\gamma^{\rm NH1}(E)\equiv 0$. This result is highly similar to that of the Hermitian Liu-Xia model, except that the energy $E$ here refers to the real part of the complex energy, which will be manifested in the subsequent IPR results. The analytical results predict that the unbound states will exist in the real-energy domain 2t<E<2t-2t<E<2t, while the bound states will exist in the real-energy domains $E>2t$ and E<2tE<-2t. For convenience, we mark $E^{\rm NH1}_{c1}=2t$ and $E^{\rm NH1}_{c2}=-2t$. Thus, the analytical boundaries should be compared with the real parts of the numerical eigenenergies.
@@
Figure 4: (Color online) The energies $E$ versus $V$ of the non-Hermitian Liu-Xia model. The colors denote the values of the IPR. The two blue solid lines denote the critical energies $E^{\rm NH1}_{c1}=2t$ and $E^{\rm NH1}_{c2}=-2t$, respectively. The energy domain $\left[E_{c2},E_{c1}\right]$ is a mixed region where unbound states and bound states coexist. Above $E_{c1}$ and below $E_{c2}$, there are bound states. The coexistence of low- and high-IPR states within the same real-energy interval signals the emergence of a mixed bound-unbound phase. The system size is $L=1597$.
@@
Figure 5: (Color online) Four wave functions of the non-Hermitian Liu-Xia model. (a) The bound state at E=45.7662tE=-45.7662t; (b) The unbound state at E=0.7628tE=-0.7628t; (c) The unbound state at E=0.7625tE=-0.7625t; (d) The bound state at $E=1.1534t$. The parameters used are $h=0.1$,$V=0.5t$, and $L=1597$.
@@
In Fig. 4, we present the real part of energies $E$ (arranged in ascending order) of the non-Hermitian Liu-Xia model as a function of the potential parameter $V$. The color coding indicates the magnitude of the IPR, and the blue lines mark the critical energies $E^{\rm NH1}_{c1}$ and $E^{\rm NH1}_{c2}$ obtained from analytical derivations. Similar to the Hermitian case, the energy region $E>E^{\rm NH1}_{c1}$ still accommodates bound states. Differently, bound states emerge in the lower real-energy domain with $E<E_{c2}$. Furthermore, it can be seen that unbound states do not vanish due to non-Hermiticity, but remain present in the energy window $E^{\rm NH1}_{c2}<E<E^{\rm NH1}_{c1}$. Different from the Hermitian systems, the conventional non-Hermitian mobility-edge and intermediate phase systems, in which unbound states fully occupy one or several entire energy domains [13, 14, 26, 20, 2, 5, 15], unbound states here appear in the form of mixed phase consisting of both unbound and bound states without a regular pattern. It is readily seen that the real part energies of unbound states are embedded in the real-part energies of bound states. This mixed phase is distinct from a conventional mobility-edge structure: the real-energy interval predicted by the Lyapunov exponent provides the possible region for unbound states, but non-Hermiticity causes localized and extended eigenstates to coexist irregularly inside this interval.
@@
Figures 5(a)-5(d) show the spatial distribution of the wave functions under four different energies. In Fig. 5(a), the real part of the energy corresponding to the wave function is located below the critical energy $E^{\rm NH1}_{c2}$. It can be observed that its spatial distribution state is localized and it is a bound state. The real-part energies of the wave functions in Figs. 5(b)-5(c) are located within the energy domain $E^{\rm NH1}_{c2}<E<E^{\rm NH1}_{c1}$. The distributions in Figs. 5(b) and 5(d) are extended, corresponding to unbound states, while the distribution in Fig. 5(c) is localized, corresponding to a bound state. This result clearly demonstrates that in non-Hermitian cases, within the energy domain $E^{\rm NH1}_{c2}<E<E^{\rm NH1}_{c1}$, bound states and unbound states coexist, which is consistent with the results of IPR. This also indicates that the analytical solution derived based on Avila’s global theory can predict both the energy domain of unbound states and the energy domain of pure bound states.
@@
Next, we study the non-Hermitian effect on the deeper potential model. Considering on-site gain and loss in $V_{2}(n)$, the resulting non-Hermitian potential $V^{\rm NH2}_{2}(n)$ is given by
$V^{\rm NH2}_{2}(n)$ $=\frac{V\left[1+\sin^{2}(2\pi\alpha n+\theta+ih)\right]}{\cos^{2}(2\pi\alpha n+\theta+ih)}$ (21)
@@
Following the previous strategy, here we again express $T^{\rm NH2}_{n}$ as the product of two parts, i.e., $T^{\rm NH2}_{n}=A^{\rm NH2}_{n}B^{\rm NH2}_{n}$, and thus we have
@@
It is noted that $A^{\rm NH1}_{n}=A^{\rm NH2}_{n}$; therefore, we have
@@
To obtain $\gamma^{B}_{\rm NH2}$, we again employ Avila’s global theory. After performing the analytic continuation ($\theta\rightarrow\theta+i\epsilon$) on the matrix $B^{\rm NH2}_{n}$, its expression in the infinite-$\epsilon$ limit reads
@@
Then, $\gamma^{B}_{\rm NH2}$ is given by